OpenStudy (anonymous):

f(x)=2(x^(1/2))-Ax. A>0. find the lim f'(x) as x goes to 0. I'm thinking that the answer is infinity, but I was just wondering if someone could please check it

OpenStudy (anonymous):

\[f(x)=2\sqrt{x}-Ax\] right?

OpenStudy (anonymous):

so \[f'(x)=\frac{1}{\sqrt{x}}-A\]

OpenStudy (anonymous):

yeah, so then the derivative is x^(-1/2)+A

OpenStudy (anonymous):

there is no limit as \(x\to 0\) or i guess you could say \(\infty\)

OpenStudy (anonymous):

oh, i meant -A

OpenStudy (anonymous):

ok, thanks

OpenStudy (anonymous):

no matter, still no limit

OpenStudy (anonymous):

I also have another question, could you check and see if the second derivative of the original function is -1/(2x(x^1/2))

OpenStudy (anonymous):

sorry i don't know how to do the square root

OpenStudy (anonymous):

\sqrt{a}

OpenStudy (anonymous):

but your answer is correct in any case

OpenStudy (anonymous):

ok, thank you!

OpenStudy (anonymous):

\[-\frac{1}{2x^{\frac{3}{2}}}\] is one version, or what you wrote yw